OpenStudy (azionne):

7(x -2) + 3(x + 2) = 5(x - 6)

OpenStudy (pawanyadav):

7x-14 +3x+6=5x-30 10x -8=5x-30 I think you can do it now

OpenStudy (alexandervonhumboldt2):

first distribute: a*(b-c)=ab-ac thus 7*(x-2)=?

pooja195 (pooja195):

Start by distributing. \[\rm~(7)(x)+(7)(−2)+(3)(x)+(3)(2)=(5)(x)+(5)(−6)\] Write the equation you get.